OpenStudy (anonymous):

need to find horizontal asymptote of (sin x)/( 1+ cos x) using limit when x->infinity

OpenStudy (anonymous):

1+cosx=0 when x=pie rad= 180 degree

OpenStudy (anonymous):

:/ im confused

OpenStudy (anonymous):

i think that would be a vertical asymptote, not a horizontal

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

wow, so it turns out that \[\frac{\sin(x)}{1+\cos(x)} = \tan(\frac{x}{2}) \] tan doesnt have any horizontal asymptotes...so i guess this function doesnt either?

OpenStudy (anonymous):

okay, thanks! :)

myininaya (myininaya):

gj joe i didn't see it was =tan(x/2)

myininaya (myininaya):

it would have been alot easier if i had seen that

myininaya (myininaya):

just so everyone knows veritcal asymptotes or holes occur when the function does not exist horizontal asymptotes are those thingys you find when you let x approach infinity

OpenStudy (anonymous):

"those thingys" lol

myininaya (myininaya):

it is my mathematical notation it is my mathematical variable

OpenStudy (anonymous):

(not that i can think of a better name at this time >.>)

myininaya (myininaya):

:)

myininaya (myininaya):

i could not sleep its 6:49 am here

OpenStudy (anonymous):

go into the thread by aross, help me solve that 10x 10 matrix lol

myininaya (myininaya):

10 by 10 thats crazy do that using an algorithm on a computer

OpenStudy (anonymous):

i put it in my phone, its actually solvable =/

myininaya (myininaya):

ok i will look at it

OpenStudy (anonymous):

another question...finding critical points of (e^-x) (sin x)

OpenStudy (anonymous):

because I know that in division, I just have to look for the first derivative, and the value that make the numerator = 0, but not the denominator, is the critical point...but in this case, which one is the critical point?

myininaya (myininaya):

f'(x)=-e^(-x)(sinx)+e^(-x)*cosx f'(x)=e^(-x){-sinx+cosx)=0 e^(-x) never 0 so we have -sinx+cosx=0 sinx=cosx sin and cos are the same when x=pi/4 and when x=5pi/4 so we can generate a thingy for all the critical numbers x=2npi+pi/4 and x=2npi+5pi/4 for n=...,-3,-2,-1,0,1,2,3,...

OpenStudy (anonymous):

sounds good to me.

myininaya (myininaya):

i added 2npi because we can keep going around the circle over and over again until we get dizzy and then some more

OpenStudy (anonymous):

the sun is coming up <.<

myininaya (myininaya):

the sun is up

myininaya (myininaya):

where is the 10 by 10 matrix?

OpenStudy (anonymous):

so the critical point of (e^-x)(sin x) are 2n pi+pi/4 and 2n pi + 5pi/4

myininaya (myininaya):

and since you say critical point we should say (2npi+pi/4,f(2npi+pi/4)) and (2npi+5pi/4,f(2npi+5pi/4))

myininaya (myininaya):

i named your function f by the way i hope that was okay

OpenStudy (anonymous):

it's okay ;)

myininaya (myininaya):

:)

OpenStudy (anonymous):

myininaya, I have to find the intervals in which f(x) is increasing or decreasing....i read that I have to look for the zeros of the first derivative of f(x)...what else do I have to do after looking for the zeroes of f'(x)??....what f'(x) doesn't have zeroes?? ....how do I find the increasing or decreasing intervals??

OpenStudy (anonymous):

for example in (x^2 -16) / (x^2 -4x)

OpenStudy (anonymous):

the first derivative is -4(x^2) but that doesn't have zeroes or roots...how do I know the increasing and decreasing intervals?

myininaya (myininaya):

ok f tells where the number is f' tells us if the function is increasing, decreasing, or neither f'' tells us if the function is concave up or concave down so we need f' to find intervals of increase and decrease and where it is neither(when the slope is 0) so if \[f(x)=\frac{x^2-16}{x^2-4x}=\frac{(x-4)(x+4)}{x(x-4)}=\frac{x+4}{x},x \neq 4=1+4x^{-1}, x \neq 4\] \[f'(x)=-4x^{-1-1}=-4x^{-2}=\frac{-4}{x^2}\] ok we x=0 is a critical number -----0---- (-inf,0) (0,inf) we choose test numbers before and after each critical number so we choose a number before 0 (doesn't matter what number) f'(-3)=a negative number => the function is decreasing on (-inf,0) now we choose a number after 0 (doesn't matter what number) f'(5)=a negative number => the function is also decreasing on (0,inf)

myininaya (myininaya):

my definition of critical number is any numbers that make f' not exist and also satisfy f'=0

myininaya (myininaya):

we do not have the latter here but the first part we do

OpenStudy (anonymous):

okay...so, a critical point or critical number is a number that makes the equation = 0?

myininaya (myininaya):

derivative of equation=0

OpenStudy (anonymous):

f'(x) = 0 or f'(x) = does not exist this are critical numbers?

myininaya (myininaya):

some people might also add a little bit to the defintion of the critical number and say: any number in the domain of f that makes f' not exist or f'=0 is a critical number

myininaya (myininaya):

yes!

myininaya (myininaya):

i do not use the domain part very much like for instance we said above f' does not exist when x=0 above so x=0 is a critical number some people might say x=0 is not a critical number because 0 is not in the domain of f so they might just choose any number from -inf to inf to see if the function is decreasing or increasing and in fact we see that f'<0 for all numbers except at x=0 (since we have discontinuity; you can also say it isn't doing anything at x=4 since f doesn't exist there)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

let me write this down...this is key for me

myininaya (myininaya):

just remember critical numbers are numbers satisfying f' does not exist and f'(x)=0 and you will be golden!

myininaya (myininaya):

i said and lets say or

OpenStudy (anonymous):

perfect!

myininaya (myininaya):

well you should be golden on finding where f is increasing/decreasing/neither

myininaya (myininaya):

of course there is alot more cal to learn and not all of it is about critical numbers

OpenStudy (anonymous):

you're right!

OpenStudy (anonymous):

inflection points are values that make the second derivative = 0 or not exist?

myininaya (myininaya):

yes! :)

myininaya (myininaya):

well let me add two more things it must also satisfy that the concavity switches at that number to be inflection point it also must be in the domain of the function to be an inflection point

OpenStudy (anonymous):

it is possible to not have inflection points??..like here (x^2 -16) / (x^2 -4x)

OpenStudy (anonymous):

if a value of the critic point isn't in the domain of the function, then that value isn't a critic point?

OpenStudy (anonymous):

or that only applies for inflection points?

myininaya (myininaya):

f''(x)=8/(x^3) f'' dne at x=0 but 0 is not in the domain of f (so there is no inflection point) we can choose test number around 0 to plug into f'' to see when f is concave up/down ---0--- f''(-1)=negative number => concave down f''(1)=positive number => concave up critical numbers i would say do not have to be in the domain of f but critical points do

myininaya (myininaya):

inflection points also have to be in the domain to be considered an inflection point

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

hi again myininaya!

OpenStudy (anonymous):

I need to graph few equations, if I don't know how to graph, I can choose random x values and substitute them in the equation?